=> x=y => x*x = y*x => x2 = yx => x2-y2 = yx- y2 => = y => x+y = y => y+y =y => 2y* 1 = y *1 => 2y / y = 1 => 2 =1 prove that this is wrong

=>  x=y
=> x*x = y*x
=>  x2 =  yx
=>  x2-y2 = yx- y2
=>   = y
=>  x+y = y
=>  y+y =y
=> 2y* 1 = y *1
=>   2y / y = 1
=>    2 =1
prove that this is wrong

Grade:9

2 Answers

purva gupta
8 years ago
x^2-y^2=yx-y^2
(x-y)(x+y)=y(x-y)
x-y=0
since 0/0 is undefined
u cant cut (x-y) on either side
and therefore writting (x+y)=y is wrong
manmath
7 years ago
 
hii friend,
x^2-y^2=yx-y^2 (x-y)(x+y)=y(x-y) x-y=0 since 0/0 is undefined u cant cut (x-y) on either side and therefore writting (x+y)=y is wrong